# Dynamically localized but deconfined excitations in strained classical spin ice

### v1

@@ IV.3 Magnetic noise
The stochastic thermal motion of monopoles generates magnetic noise, which is experimentally observable and has been measured in spin ice materials [11, 12]. Within the spin ice regime $-1<\delta<1$, we characterize the thermal magnetic noise of our model by extracting the power spectral density (PSD) for the different spin subspecies. The subspecies PSD is defined as
@@
Figure 8: Power spectral density (PSD) of the magnetic noise, cf. Eq. (missing) 9, measured from spin subspecies $\mathcal{S}_{[111]}$ (top row( and $\mathcal{S}_{K}$ (bottom row). The frequency is in units of the spin-flip rate, which is, e.g., for Dy2Ti2O7 approximately $10^{4}\ {\rm s}^{-1}$. We show data for the different spin ice regimes $-1<\delta\leq 0$ and $10\leq\delta<$ 1 in the left and right columns, respectively. The values for $\delta$ and $T$ are chosen such that a finite but small number of cheap monopoles are present in our simulation of $N=4000$ spins. For $\delta<0$, type 1A monopoles are energetically favored and propagate diffusively, yielding a Lorentzian noise spectrum for both subspecies in the left column, panel (a,c). Motion of type 1B monopoles, which are favored for $\delta>0$, are much more constrained and are displayed in the right column. Since type 1B monopoles cannot propagate through $\mathcal{S}_{[111]}$ spins, the Lorentzian observed in panel (b) is attributed to the creation, annihilation, and propagation of expensive monopoles. Thus, on times scales of expensive type 1A monopoles, $\kappa_{1}$ from Eq. (missing) 8, we find a perfect scaling collapse in the inset when considering rescaled units. The anisotropy of our model manifests itself most prominently in panel (d), revealing a suppressed noise for intermediate frequencies coming from the localization of cheap monopoles.
@@
Focusing on type 1A dynamics first ($-1<\delta\leq 0$; left column of Fig. 8), we find that the PSDs of all spin subspecies are essentially Lorentzian, consistent with the predicted anisotropic diffusive monopole motion observed in panel (a) of Fig. 6. The PSD of subspecies $\mathcal{S}_{[111]}$ depends only weakly on $\delta$, which emphasizes the fact that type 1A monopoles are always free to move along $[111]$. The PSD of sublattice $\mathcal{S}_{K}$ displays a stronger $\delta$-dependence, and for increasing $\delta$, the low-frequency plateau appears to shift to lower frequency and larger amplitude, consistent with an increase in the relaxation time $\tau$. This is a manifestation of the increased energy cost associated with converting a type 1A to a type 1B monopole, which reduces the monopole mobility through subspecies $\mathcal{S}_{K}$.
Type 1B dynamics
@@
The monopole motion is strongly constrained for $\delta>0$, and this manifests itself in the PSD ($10\leq\delta<$ 1; right column of Fig. 8). The fluctuations of spins belonging to $\mathcal{S}_{[111]}$, panel (b), come purely from the motion, creation, or annihilation of type 1A monopoles, which are now energetically more expensive than type 1B. The resulting PSD is Lorentzian, as in the $\delta<0$ case but with significantly suppressed magnitude and an increased relaxation time owing to the lower density of type 1A monopoles in this regime. Rescaling the PSD and frequency by $\kappa_{1}$ from Eq. (missing) 8, proportional to the density to type 1A monopoles, we find a perfect scaling collapse as illustrated in the inset of panel (b) in Fig. 2; see also subsection B.2 for further details.
@@
The PSD for spins belonging to $\mathcal{S}_{K}$, panel (d) of Fig. 8, behaves quite differently in the $\delta>0$ regime. Here, the high-frequency noise is primarily generated by the localized motion of type 1B monopoles. For $\omega\gtrsim 1/\tau_{0}$ the PSD displays a $\omega^{-2}$ decay, which is typical for PSD curves measured in Monte Carlo simulations [12]. For larger values of $\delta$, there is a regime of intermediate frequencies where the PSD decays significantly slower with frequency – a manifestation of the monopole localization. This slower decay would turn into a plateau persisting to arbitrarily small frequencies as $T\to 0$ if the only relaxation mechanism in the system was the zero-energy motion of type 1B monopoles. However, both the presence of a small but non-zero density of type 1A monopoles and the spontaneous creation of new type 1B monopoles offer alternate avenues for the system to relax. This manifests as a faster decay of the noise spectrum at low frequencies. Below some frequency, the noise must plateau in all cases (unless $T\to 0$) as all spins can flip at a finite energy cost, and beyond some time scale, the system is therefore completely uncorrelated from its previous configurations. This frequency is, however, not accessible in our simulations for $\delta>0$.
@@
Experimentally, there has recently been renewed interest in the dynamics of magnetic monopoles [11, 33, 12, 10, 34], largely driven by the ability to probe magnetic noise spectra using SQUIDs. The magnetic noise plots in Sec. IV took advantage of the microscopic nature of the Monte Carlo simulation to present sublattice-resolved PSDs. For completeness, we present data for the (experimentally accessible) directional PSD curves in Fig. 9. These can be obtained from the sublattice resolved data in Fig. 8. We choose moderate values of $\delta/T$ that might be realized in future experiments using the sublattice doping discussed in the next section. As expected, we find stark differences between the $[111]$ direction and the orthogonal direction $[1\bar{1}0]$. Fig. 9 also demonstrates the transition between short-time anisotropic and long-time isotropic behavior. This crossover occurs on the time scale $\sim\kappa_{0}$, beyond which the distinction between the monopole types loses its meaning.
@@ V Experimental realizations
Uniaxial pressure is more limited than hydrostatic pressure, with previous experiments achieving pressures of 2.22.22.2\,2.2GPa (1.31.31.3\,1.3GPa) for Ho2Ti2O7 (Dy2Ti2O7[15, 16, 17]. By modeling the magnetization with respect to the applied pressure and an external field, Refs. [16, 17] estimate the increase in the nearest-neighbor coupling parallel to the pressure axis and the lattice compression. Taking both, dipolar and exchange interactions with the dipolar spin ice Hamiltonian [44], into account, we expect this to correspond to an increase of the effective nearest-neighbor coupling by $\delta\sim 0.06$ and $\delta\sim 0.02$ for Ho2Ti2O7 and Dy2Ti2O7, respectively.
@@
Another mechanism to achieve larger magnitudes of $|\delta|$ involves substituting spins located on the sublattice $\mathcal{S}_{[111]}$ with a species of spins characterized by a different magnetic moment. This approach might be capable of exploring the full phase diagram in Fig. 2, which contains ferro- and antiferromagnetic bonds. For example, introducing spins with a larger or smaller magnetic moment, like doping Ho atoms into the $\mathcal{S}_{[111]}$ positions within Dy2Ti2O7, naturally induces variations in $\delta$.
Extending this idea, atoms that interact antiferromagnetically with the other species open up the possibility to observe phases predicted for the $\delta<-1$. A notable example of this approach are tripod kagome systems where atoms on the sublattice $\mathcal{S}_{[111]}$ are substituted with non-magnetic ions [7, 8]. They realize the extremal case $\delta=-1$, where we have disconnected kagome planes. Note, however, that doping atoms at the right position is extremely challenging, and it is much easier to place them randomly within the crystalline structure. Hence, understanding and modeling the Hamiltonian where $\delta$ is randomly assigned to a subset of spins is potentially important future work.
Lastly, a growing effort in artificial spin ice systems [45] might provide an avenue to realize our model and confirm the relation between Monte Carlo simulation and real-time dynamics.
@@ VI Conclusions
Our simple model Hamiltonian $H_{S}$ in Eq. (missing) 2, nominally describing spin ice under uniaxial $[111]$-strain, has turned out to support a rich and varied physical landscape. Using the grouping of the sixteen single-tetrahedron configurations into four degenerate types, we have successfully described the full phase diagram of $H_{S}$. Remarkably, the extensive degeneracy of the ice ground states is preserved for a wide range of distortion values, which is in stark contrast to the behavior of uniaxial pressure along other crystallographic axes [19]. In contrast, the thermodynamic properties contain information about the distortion, for example, in the form of two peaks in the specific heat. The distortion further manifests itself in the dynamical properties, with two dynamical regimes appearing in the low-temperature limit. Using single-spin flip dynamics, we find that the magnetic monopoles split into a diffusive and a localized type, which dominate the dynamics at negative and positive $\delta$, respectively. For $\delta>0$, the localization of the low-energy monopoles suppresses the magnetic noise in a characteristic, anisotropic fashion. The time scale on which the monopoles remain localized grows exponentially with inverse temperature, making their localization a relevant property at low temperatures regardless of the strength of the distortion. This, presumably, also leads to an exponential slow-down of thermalization at low temperatures.
Including additional dynamic processes, such as the ring exchange updates discussed here, partly restores the diffusive motion of the localized monopoles. In this case, the monopoles remain localized in the strain direction – effectively making them sub-dimensional excitations.
@@ Acknowledgement
Appendix A Numerical methods
@@
For thermodynamic measurements, we simulated the system by annealing from a high temperature, applying both worm updates and single-spin flip updates at each temperature step. As mentioned above, we only performed simulations for $\delta\geq-1$ and then invoke the duality given in Eq. (missing) 4 to deduce the behavior for $\delta<-1$.
The equilibrium system dynamics were simulated using the standard model of spin ice dynamics [28, 9]: Metropolis algorithm Monte Carlo using only single spin flip updates and defining the flipping attempt time $\tau_{0}=1$ MC-sweep (1 MC-sweep corresponds to $N$ attempted spin flips). A combination of single-spin flip and worm updates were used to initialize the system at the desired temperature, and the dynamical measurements were only begun once the system was in thermal equilibrium. The PSD was computed from time traces of the sublattice magnetization using Welch’s method [46]. In an infinite system, the monopole density only goes to zero as temperature goes to zero. However, in a finite system, if the temperature is chosen too low, all monopoles can vanish, preventing any magnetization dynamics. Thus, when measuring the magnetization noise, we ensured that we were always at temperatures where the actual monopole number did not drop to zero at any point during the simulation.
Finally, the MSDs of individual monopoles were measured by first cooling a system of 54,000 spins until an ice state without any monopoles was reached. A single pair of low-energy monopoles is then created at an arbitrary position. One of these monopoles was kept fixed at its original position, whereas the other monopole was allowed to move. Each Monte Carlo update now consists of randomly selecting one of the four spins surrounding the mobile monopole and flipping this with probability $\min[1,\exp{(-\Delta E/T}]$ if the selected spin is a majority spin. As creation and annihilation are not allowed, the monopole dynamics only depend on the ratio $J\delta/T$. We define four such updates to be equivalent to time $\tau_{0}$. This mimics the motion of a monopole under standard model dynamics in the limit of very low monopole density. To minimize the impact of the static monopole on the measurement results, the mobile monopole was allowed to move for approximately $10^{6}\tau_{0}$ before its trajectory was measured. The MSD curves shown in Fig. 6 and Fig. 12 are computed by performing a bootstrap average of the trajectories of 200 independent monopoles.
@@
Appendix B Further numerical results
B.1 Phase transition at $\delta>1$
We have identified a symmetry-breaking phase transition upon cooling when $\delta>1$ (and $\delta<-3$), cf. see Sec. II and III. This is a conventional paramagnet to ferromagnet phase transition in an Ising spin system, and we expect it to fall within the 3d Ising universality class [47, 19]. In the thermodynamic limit, one would then expect the susceptibility to diverge as $\chi\sim\left|T-T_{c}\right|^{-\gamma}$ and the specific heat to diverge as $C_{v}\sim\left|T-T_{c}\right|^{-\alpha}$ as $T$ approaches $T_{c}$ from above. In Fig. 10, we plot the specific heat and the susceptibility for all spin subspecies for temperatures close to the critical temperature, approaching $T_{c}$ from above. The critical temperature is taken to be the temperature at which we observe the maximal specific heat and susceptibility value. We display results for two different linear system sizes, $L=6$ and $10$, and find that the growth of both the susceptibility and specific heat approximately agree with the predicted critical exponents: $\gamma=1.23$ and $\alpha=0.11$ [48]. This is consistent with the thermal phase transitions being 3d Ising transitions, as expected.

### v2

@@ IV.3 Magnetic noise
The stochastic thermal motion of monopoles generates magnetic noise, which is experimentally observable and has been measured in spin ice materials [11, 12, 34, 35]. Within the spin ice regime $-1<\delta<1$, we characterize the thermal magnetic noise of our model by extracting the power spectral density (PSD) for the different spin subspecies. The subspecies PSD is defined as
@@
Figure 8: Power spectral density (PSD) of the magnetic noise, cf. Eq. (missing) 10, measured from spin subspecies $\mathcal{S}_{[111]}$ (top row) and $\mathcal{S}_{k}$ (bottom row). The frequency is in units of the spin-flip rate, which is, e.g., for Dy2Ti2O7 approximately $10^{4}\ {\rm s}^{-1}$. We show data for the different spin ice regimes $-1<\delta\leq 0$ and $10\leq\delta<$ 1 in the left and right columns, respectively. The values for $\delta$ and $T$ are chosen such that a finite but small number of cheap monopoles are present in our simulation of $N=4000$ spins. For $\delta<0$, type 1A monopoles are energetically favored and propagate diffusively, yielding a Lorentzian noise spectrum for both subspecies in the left column, panel (a,c). Motion of type 1B monopoles, which are favored for $\delta>0$, are much more constrained and are displayed in the right column. Since type 1B monopoles cannot propagate through $\mathcal{S}_{[111]}$ spins, the Lorentzian observed in panel (b) is attributed to the creation, annihilation, and propagation of expensive monopoles. Thus, on times scales of expensive type 1A monopoles, $\kappa_{1}$ from Eq. (missing) 9, we find a perfect scaling collapse in the inset when considering rescaled units. The anisotropy of our model manifests itself most prominently in panel (d), revealing a suppressed noise for intermediate frequencies coming from the localization of cheap monopoles.
@@
Focusing on type 1A dynamics first ($-1<\delta\leq 0$; left column of Fig. 8), we find that the PSDs of all spin subspecies are essentially Lorentzian, consistent with the predicted anisotropic diffusive monopole motion observed in panel (a) of Fig. 6. The PSD of subspecies $\mathcal{S}_{[111]}$ depends only weakly on $\delta$, which emphasizes the fact that type 1A monopoles are always free to move through the $\mathcal{S}_{[111]}$ spins. The PSD of sublattice $\mathcal{S}_{k}$ displays a stronger $\delta$-dependence, and for increasing $\delta$, the low-frequency plateau appears to shift to lower frequency and larger amplitude, consistent with an increase in the relaxation time $\tau$. This is a manifestation of the increased energy cost associated with converting a type 1A to a type 1B monopole, which reduces the monopole mobility through subspecies $\mathcal{S}_{k}$.
Type 1B dynamics.
@@
The monopole motion is strongly constrained for $\delta>0$, and this manifests itself in the PSD ($10\leq\delta<$ 1; right column of Fig. 8). The fluctuations of spins belonging to $\mathcal{S}_{[111]}$, panel (b), come purely from the motion, creation, or annihilation of type 1A monopoles, which are now energetically more expensive than type 1B. The resulting PSD is Lorentzian, as in the $\delta<0$ case, but with significantly suppressed magnitude and an increased relaxation time owing to the lower density of type 1A monopoles in this regime. Rescaling the PSD and frequency by $\kappa_{1}$ from Eq. (missing) 9, proportional to the density to type 1A monopoles, we find a perfect scaling collapse as illustrated in the inset of panel (b) in Fig. 8; see also subsection C.2 for further details.
@@
The PSD for spins belonging to $\mathcal{S}_{k}$, panel (d) of Fig. 8, behaves quite differently in the $\delta>0$ regime. Here, the high-frequency noise is primarily generated by the localized motion of type 1B monopoles. For $\omega\gtrsim 1/\tau_{0}$ the PSD displays a $\omega^{-2}$ decay, which is typical for PSD curves measured in Monte Carlo simulations [12]. For larger values of $\delta$, there is a regime of intermediate frequencies where the PSD decays significantly slower with frequency – a manifestation of the monopole localization. This slower decay would turn into a plateau persisting to arbitrarily small frequencies as $T\to 0$ if the only relaxation mechanism in the system was the zero-energy motion of type 1B monopoles. However, both the presence of a small but nonzero density of type 1A monopoles and the spontaneous creation of new type 1B monopoles offer alternate avenues for the system to relax. This manifests as a faster decay of the noise spectrum at low frequencies. Below some frequency, the noise must plateau in all cases (unless $T\to 0$) as all spins can flip at a finite energy cost, and beyond some timescale, the system is therefore completely uncorrelated from its previous configurations. This frequency is, however, not accessible in our simulations for $\delta>0$.
@@
Experimentally, there has recently been renewed interest in the dynamics of magnetic monopoles [11, 36, 12, 10, 37], largely driven by the ability to probe magnetic noise spectra using SQUIDs. The magnetic noise plots in Sec. IV took advantage of the microscopic nature of the Monte Carlo simulation to present sublattice-resolved PSDs. For completeness, we present data for the (experimentally accessible) directional PSD curves in Fig. 9. These can be obtained from the sublattice resolved data in Fig. 8. We choose moderate values of $\delta/T$ that might be realized in future experiments using the sublattice doping discussed in the next section. As expected, we find stark differences between magnetic noise measured along the $[111]$ direction and along its orthogonal direction $[1\bar{1}0]$. Fig. 9 also demonstrates the transition between short-time anisotropic and long-time isotropic behavior. This crossover occurs on the timescale $\sim\kappa_{0}$, beyond which the distinction between the monopole types loses its meaning.
@@ V Experimental realizations
Uniaxial stress is more limited than hydrostatic pressure, with previous experiments demonstrating stresses of up to 2.22.22.2\,2.2GPa (1.31.31.3\,1.3GPa) for Ho2Ti2O7 (Dy2Ti2O7[15, 16, 17, 18]. By modeling the magnetization with respect to the applied stress and an external field, Refs. [16, 17] estimate the increase in the nearest-neighbor coupling parallel to the stress axis and the lattice compression. Taking both dipolar and exchange interactions, with the dipolar spin ice Hamiltonian [46], into account, we expect this to correspond to an increase of the effective nearest-neighbor coupling of $\delta\sim 0.06$ and $\delta\sim 0.02$ for Ho2Ti2O7 and Dy2Ti2O7, respectively.
@@
Another mechanism to achieve larger magnitudes of $|\delta|$ involves substituting spins located on the sublattice $\mathcal{S}_{[111]}$ with a species of spins characterized by a different magnetic moment. This approach might be capable of exploring the full phase diagram in Fig. 2, which contains ferromagnetic and antiferromagnetic bonds. For example, introducing spins with a larger or smaller magnetic moment into the $\mathcal{S}_{[111]}$ positions of a spin compound leads to variations in $\delta$.
Extending this idea, atoms that interact antiferromagnetically with the other species open up the possibility to observe phases predicted for the $\delta<-1$. A notable example of this approach are tripod kagome systems where atoms on the sublattice $\mathcal{S}_{[111]}$ are substituted with nonmagnetic ions [7, 8]. They realize the extremal case $\delta=-1$, where we have disconnected kagome planes. Note, however, that doping atoms at the right position is extremely challenging, and it is much easier to place them randomly within the crystalline structure. Hence, understanding and modeling the Hamiltonian where $\delta$ is randomly assigned to a subset of spins is potentially important future work.
We used standard model dynamics in Sec. IV, which has proven to be an excellent model for the study of real materials [9]. It should, however, be noted that more detailed and material-specific model dynamics have been developed [47, 10]. In particular, the beyond the standard model, which replaces the single global flipping attempt rate with two rates based on a microscopic quantum description of the system, can reproduce dynamical phenomena observed in Dy2Ti2O7 (without strain) which cannot be reproduced using the standard model [10, 37]. To what extent the different model dynamics apply to specific materials under strain (or with substitutional doping) remains to be ascertained and requires quantum mechanical microscopic considerations beyond the scope of this work.
Lastly, a growing effort in artificial spin ice systems [48, 49] and thin films [50, 51, 52, 53] might provide an avenue to engineer strained models that potentially host similar localization physics. In particular, thin films allow for a variety of different strain patterns that potentially lock monopoles within the layers.
@@ VI Conclusions
Our simple model Hamiltonian $H_{S}$ in Eq. (missing) 3, nominally describing spin ice under uniaxial $[111]$-strain, has turned out to support a rich and varied physical landscape. Using the grouping of the sixteen single-tetrahedron configurations into four degenerate types, we have successfully described the full phase diagram of $H_{S}$. Remarkably, the extensive degeneracy of the ice ground states is preserved for a wide range of distortion values, which is in stark contrast to the behavior of uniaxial stress along other crystallographic axes [19]. In contrast, the thermodynamic properties contain information about the distortion, for example, in the form of two peaks in the specific heat. The distortion further manifests itself in the dynamical properties, with two dynamical regimes appearing in the low-temperature limit. Using single-spin flip dynamics, we find that the magnetic monopoles split into a diffusive and a localized type, which dominate the dynamics at negative and positive $\delta$, respectively. For $\delta>0$, the localization of the low-energy monopoles suppresses the magnetic noise in a characteristic, anisotropic fashion. The timescale on which the monopoles remain localized grows exponentially with inverse temperature, making their localization a relevant property at low temperatures regardless of the strength of the distortion. This, presumably, also leads to an exponential slow-down of thermalization at low temperatures.
Including additional dynamic processes, such as the ring exchange updates discussed here, partly restores the diffusive motion of the localized monopoles. In this case, the monopoles remain localized in the strain direction – effectively making them subdimensional excitations.
@@ Acknowledgements
Appendix A Spin and magnetization convention
Spin convention
The spins in our model are represented by classical vectors $\vec{s}$ that are constrained to point along the local $\langle 111\rangle$ directions, i.e., a spin on sublattice $\alpha$ points along the unit vector $\vec{e}_{\alpha}$ with
$\vec{e}_{1}=\frac{1}{\sqrt{3}}\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$ $\vec{e}_{2}=\frac{1}{\sqrt{3}}\begin{pmatrix}1\\ -1\\ -1\end{pmatrix}$ (11)
$\vec{e}_{3}=\frac{1}{\sqrt{3}}\begin{pmatrix}-1\\ 1\\ -1\end{pmatrix}$ $\vec{e}_{4}=\frac{1}{\sqrt{3}}\begin{pmatrix}-1\\ -1\\ 1\end{pmatrix}\,,$
where $\vec{e}_{1}$ points along the global $[111]$ direction and spins on sublattice $\mathcal{S}_{[111]}$ are aligned with $\vec{e}_{1}$. Each spin can be expressed in terms of Ising variables as $\vec{s}_{i}=\sigma_{i}\vec{e}_{\alpha}$ where we take the spin length to be one. In our convention, $\sigma=+1$ corresponds to a spin pointing from an “up” tetrahedron to a “down” tetrahedron. With respect to Fig. 1, this refers to pointing from a blue to a red tetrahedron.
When rewriting our Hamiltonian, Eq. (missing) 1, in terms of the Ising variables we make use of the fact that $\vec{s}_{i}\cdot\vec{s}_{j}=\sigma_{i}\sigma_{j}\vec{e}_{\alpha}\cdot\vec{e}_{\beta}=-\sigma_{i}\sigma_{j}/3$ for any nearest-neighbor pair of spins.
Magnetization convention
Two different versions of the magnetization are used in this work. The directional magnetization is defined as
$M_{\vec{n}}=\vec{n}\cdot\sum_{i}\vec{s}_{i}\,,$ (12)
for some unit vector $\vec{n}$. The equilibrium $[111]$-directional susceptibility displayed in Fig. 4 is given by
$\chi_{[111]}=\frac{1}{N}\left.\frac{dM_{[111]}}{dH}\right|_{H=0}=\frac{1}{N}\frac{\left.\langle M_{[111]}^{2}\right\rangle-\left\langle M_{[111]}\right\rangle^{2}}{k_{B}T}\,.$ (13)
In contrast, the sublattice magnetization is defined as
$M_{\cal S}=\sum_{i\in\mathcal{S}}\sigma_{i}\,.$ (14)
where the sum is over all sites on sublattice ${\cal S}$. The magnetic noise computed both from the directional and the sublattice magnetization are discussed in Sec. IV.
Appendix B Numerical methods
@@
For thermodynamic measurements, we simulated the system by annealing from a high temperature, applying both worm updates and single-spin flip updates at each temperature step. As mentioned above, we only performed simulations for $\delta\geq-1$ and then invoke the duality given in Eq. (missing) 5 to deduce the behavior for $\delta<-1$.
The equilibrium system dynamics were simulated using the standard model of spin ice dynamics [29, 9]: Metropolis algorithm Monte Carlo using only single spin flip updates and defining the flipping attempt time $\tau_{0}=1$ MC-sweep (1 MC-sweep corresponds to $N$ attempted spin flips). A combination of single-spin flip and worm updates are used to initialize the system at the desired temperature, and the dynamical measurements were only begun once the system was in thermal equilibrium. The PSD was computed from time traces of the magnetization using Welch’s method [54]. In an infinite system, the monopole density only goes to zero as temperature goes to zero. However, in a finite system, if the temperature is chosen too low, all monopoles can vanish, preventing any magnetization dynamics. Thus, when measuring the magnetization noise, we ensured that we were always at temperatures where the actual monopole number did not drop to zero at any point during the simulation.
Finally, the MSDs of individual monopoles were measured by first cooling a system of 54,000 spins until an ice state without any monopoles was reached. A single pair of low-energy monopoles is then created at an arbitrary position. One of these monopoles was kept fixed at its original position, whereas the other monopole was allowed to move. Each Monte Carlo update now consists of randomly selecting one of the four spins surrounding the mobile monopole and flipping this with probability $\min[1,\exp(-\Delta E/T)]$ if the selected spin is a majority spin. As creation and annihilation are not allowed, the monopole dynamics only depend on the ratio $J\delta/T$. We define four such updates to be equivalent to time $\tau_{0}$. This mimics the motion of a monopole under standard model dynamics in the limit of very low monopole density. To minimize the impact of the static monopole on the measurement results, the mobile monopole was allowed to move for approximately $10^{6}\tau_{0}$ before its trajectory was measured. The MSD curves shown in Fig. 6 and Fig. 12 are computed by performing a bootstrap average of the trajectories of 200 independent monopoles.
@@
Appendix C Further numerical results
C.1 Phase transition at $\delta>1$
We have identified a symmetry-breaking phase transition upon cooling when $\delta>1$ (and $\delta<-3$), cf. see Sec. II and III. This is a conventional paramagnet to ferromagnet phase transition in an Ising spin system, and we expect it to fall within the 3d Ising universality class [55, 19]. In the thermodynamic limit, one would then expect the susceptibility to diverge as $\chi\sim\left|T-T_{c}\right|^{-\gamma}$ and the specific heat to diverge as $C_{v}\sim\left|T-T_{c}\right|^{-\alpha}$ as $T$ approaches $T_{c}$ from above.
In Fig. 10, we plot the specific heat and the susceptibility for all spin subspecies for temperatures close to the critical temperature, approaching $T_{c}$ from above. The critical temperature is taken to be the temperature at which we observe the maximal specific heat and susceptibility value. We display results for two different linear system sizes, $L=6$ and $10$, and find that the growth of both the susceptibility and specific heat approximately agree with the predicted critical exponents: $\gamma=1.23$ and $\alpha=0.11$ [56]. This is consistent with the thermal phase transitions being 3d Ising transitions, as expected.